A stationary car on a train track, m1 = 1500 kg, is struck by a train, m2 = 4500 kg. After the collision, the train and the car move together with a velocity of +2.3 m/s. Assume no external forces act during the collision.

a) What type of collision is this, and what physical quantity is conserved?

b) Calculate the initial velocity of the train in km/h.

c) Calculate the change in kinetic energy of the system.
Answer #1

a) This is an inelastic collision and momentum is conserved in such collisions.

b) Let u be the initial velocity of the train.

Then,

4500*u = (1500+4500)*2.3

=> u = 3.07 m/s

c) Hence, the initial KE of the system is (1/2)m2u^2 + 0 = (1/2)*4500*3.07^2 = 21206 J

Final KE = (1/2)(4500+1500)(2.3)^2 = 15870 J

Hence change in KE = 21206 - 15870 = 5336J decrease